OpenStudy (anonymous):

find the number of real solutions for the equlation 3x^2+6x+3=0

OpenStudy (anonymous):

\[\Delta=6^{2}-4\times3\times3=0\] so there is only one solution

OpenStudy (anonymous):

\[3x^2+6x+3=0\] \[3(x^2+2x+1)=0\] \[3(x+1)^2=0\] \[x=-1\]